A company is considering a three-year project. New equipment will cost $200,000. The equipment falls in the MACRS three-year class (.3333, .4445, .1481, .0741) it will have a salvage value at the end of the Project of $50,000. The project is expected to produce sales of $100,000 in the first year and sales will increase by $50,000 each year after that. Expenses are expected to be 40% of sales. An investment in net-working capital of $5000 is required at the beginning of the project. In year 2, net working capital must be increased by $10,000. All net-working capital will be recovered in the final year. The tax rate is 30%. The company is funded with 40% debt with a yield of 7% and 60% equity with a required return of 15%.

A. What are the annual depreciation expenses and the after-tax salvage value of the machine?

B. What is the OCF for each of the three years?

C. What are the total cash flows for years 0-3?

D. What is the company's hurdle rate?

E. What is the NPV? IRR? Profitability Index? Show work.

F. Should the company invest in the project?
Answer #1

Part-A:

Depreciation Calculation:

Year Cost of Equipment Depn Rate Depreciation Closing Book Value
1                        200,000 0.3333 66660                         133,340
2                        200,000 0.4445 88900                           44,440
3                        200,000 0.1481 29620                           14,820

Book Value of Equipment at the end of project (i.e. at the end of year-3) = $14,820

Salvage value of Equipment at the end of project (i.e. at the end of year-3) = $50,000

Profit on sale of equipment = 50,000 – 14,820 = $35,180

Tax on profit on sale of equipment = 35180*30% = $10,554

After tax salvage value of the machine = 35180-10554 = $24,626.

.

.

.

Part-B:

Operating Cash Flows:

Year Sales (1) Expenses (2) = (1)*40% Depreciation (3) Operating profit (4) = (1)-(2)-(3) Tax@30% (5) = (4)*30% Operating profit after tax (6) = (4) - (5) Operating Cash Flows (7) = (6) + (3)
1                        100,000                                   40,000                                   66,660                                                     (6,660)                                   (1,998)                                                                   (4,662)                                                           61,998
2                        150,000                                   60,000                                   88,900                                                        1,100                                         330                                                                         770                                                           89,670
3                        200,000                                   80,000                                   29,620                                                     90,380                                   27,114                                                                   63,266                                                           92,886

.

Part-C:

Total Cash Flows:

Year Iniial Investment (1) Operating CashFlows(Calculated in part-b) (2) Net Working Capital (3) Salvage Value (4) Total Cash Flows (5) = (1)+(2)+(3)+(4)
0 -200000 0 -5000 0 -205000
1 0 61998 0 0 61998
2 0 89670 -10000 0 79670
3 0 92886 15000 24626 132512

.

Part-D:

Company's Hurdle Rate (i.e. Weighted Average Cost of Capital):

After tax cost of Debt = 7%*0.30 = 4.90%

Particulars Cost Ratio/Weight WACC/Hurdle rate = (Cost*Ratio/Weight)
Equity 15% 0.6 9.00%
Debt 4.90% 0.4 1.96%
WACC/Hurdle rate 10.96%

.

Part-E:

Calculation of NPV:

Year Total Cash Flows (Calculated from part-C) DF Working Discounting Factor @ 10.96% Present Value (TCF*DF)
0                                                                      (205,000) 1 1                        (205,000.00)
1                                                                          61,998 1/1.1096^1 0.901225667                             55,874.19
2                                                                          79,670 1/1.1096^2 0.812207703                             64,708.59
3                                                                        132,512 1/1.1096^3 0.731982429                             96,996.46
NPV                            12,579.23

.

IRR Calculation:

At IRR,

Present Value of Cash Out Flows = Present Value of Cash Inflows

205,000 = 61998/(1+IRR)^1 + 79670/(1+IRR)^2 + 132512/(1+IRR)^3

Computing for IRR,

we get IRR = 14.03%

.

Therefore, IRR of the project = 14.03%

.

.

Profitability Index (PI) = Present Value of Cash Inflows/Initial Investment

PI = (55874.19+64708.59+96996.46)/205000

PI = 217,579.23/205,000

PI = 1.06

.

.

Part-F:

Yes , the company should invest in the project since :

1)NPV is positive

2)IRR of he project is greater than the hurdle rate of the company (14.03%>10.96%)

3)PI of the project is greater than 1.